OpenStudy (anonymous):

Use a trigonometric substitution involving tangent to write √(961 + x^2) as a function of theta in the first quadrant without a radical So confused pleaseeee help

OpenStudy (anonymous):

would it help to know that \(\sqrt{961}=31\)?

OpenStudy (anonymous):

put \(x=32\tan(\theta)\) so \(x^2=961\tan^2(\theta)\) and you get \[\sqrt{961+x^2}=\sqrt{961+961\tan^2(\theta)}=\sqrt{961(1+\tan^2(\theta)}=31\sqrt{1+\tan^2(\theta)}\]

OpenStudy (anonymous):

sorry i meant \[x=31\tan(\theta)\]

OpenStudy (anonymous):

and now a miracle occurs, namely that \[1+\tan^2(\theta)=\sec^2(\theta)\] so you can get rid of the radical

OpenStudy (anonymous):

when you see \[\sqrt{a^2+x^2}\] use \[x=a\tan(\theta)\] and when you see \[\sqrt{a^2-x^2}\] use \[x=a\sin(\theta)\]

OpenStudy (anonymous):

we can do another if you like